The alpha value for a drug in the blood is 0.90, equating to 0.55 ng/mL. What is the concentration of total drug in the blood? The correct answer is 0.61 ng/mL. Please show how this question is solved.
Answer #1

Answer:

Step 1:

alpha = Concentration unbound / Concentration unbound + concentration bound

alpha = 0.90

Concentration unbound = 0.55ng/mL

Concentration bound = X

Step 2:

Calculate concentration bound

0.90 = 0.55 / 0.55 + X

By cross multiplying,

0.90(0.55 + X) = 0.55

(0.90 x 0.55) + (0.90 x X) = 0.55

0.495 + 0.90X = 0.55

0.90X = 0.55 - 0.495

0.90X = 0..055

X = 0.055/ 0.90

X = 0.0611

Concentration bound = 0.0611

Step 3:

Calculate the conentration of total drug in blood

Total drug = Concentration unbound + Concentration bound

= 0.55 + 0.0611

Total drug = 0.611ng/mL.

The total drug concentration in the blood = 0.61ng/mL.